OpenStudy (anonymous):

Molly brought two heads of cabbage for $1.80 . How many heads of cabbage can Willie buy if he had $28.80

OpenStudy (anonymous):

8 heads of cabbage

OpenStudy (anonymous):

Thanks Guys . :-)

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

actually, i think it's 32 because 1.80/2=28.80/x so 2(28.80)=1.80x and 57.6=1.80x. therefore 57.6/1.80=32, meaning he can buy 32 heads of cabbage.

OpenStudy (anonymous):

That's what I was getting but I needed help so Thank you .